OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. G(4, 6), H(1, 5)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

(6-5)/(4-1)=1/3 y=1/3x+14/3

OpenStudy (ajspeller):